(a) Calculate the electric potential of Q= 25×10^-9 at 1m


(b) Calculate the electric potential of Q= 25×10^-9 at 2m

1
Expert's answer
2022-02-09T06:26:15-0500

The electric potential of point charge is given by

"V=k\frac{q}{R}"

(a)

"V=9*10^9*\frac{25*10^{-9}}{1}=225\:\rm V"

(b)

"V=9*10^9*\frac{25*10^{-9}}{2}=112.5\:\rm V"